OpenStudy (ffuentes6596):

Choose the correct simplification of the expression a to the 6th power times b to the 3rd power all over a to the 4th power times b to the second power . a to the second power over b a10b5 a to the 10th power over b to the 5th power a2b (Medal for first answer)

OpenStudy (steve816):

\[\LARGE \frac{ a^6b^3 }{ a^4b^2 }=a^2b\]